Find the integrating factor to transform the given differential equation into the equation in exact differentials: (x2-3y2)dx+2xydy=0 μ=μ(x)


1
Expert's answer
2021-11-29T16:23:00-0500
"\dfrac{M_y-N_x}{N}=\dfrac{-6y-2y}{2xy}=-\dfrac{4}{x}"

Integrating factor


"\mu=\mu(x)=e^{\int(-4/x)dx}=x^{-4}"

"x^{-4}(x^2-3y^2)dx+2x^{-4}xydy=0"

"(x^{-2}-3x^{-4}y^2)dx+2x^{-3}ydy=0"

"\dfrac{\partial M}{\partial y}=-6x^{-4}y"

"\dfrac{\partial N}{\partial x}=-6x^{-4}y"

The differential equation


"(x^{-2}-3x^{-4}y^2)dx+2x^{-3}ydy=0"

is the equation in exact differentials.

Integrating factor


"\mu=\mu(x)=x^{-4}"